OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations. x – y + 2z = –2 2x + 2y + z = 7 3x + 3y – z = 3

OpenStudy (blurbendy):

to eliminate z in the first and third equations, you have to multiply the entire third equation by 2

OpenStudy (blurbendy):

then add them together

OpenStudy (anonymous):

so it would be 6x+6y-2z=6 then what do you mean by add them together?

OpenStudy (blurbendy):

x - y +2x = 2 + 6x + 6y -2z = 6 = 7x +5y = 8 z's cancel

OpenStudy (blurbendy):

now, to eliminate z in the second and third equations, just add them together

OpenStudy (blurbendy):

that shouldn't be a 2x in the first equation, it should be 2z

OpenStudy (blurbendy):

wrote 2x by accident ><

OpenStudy (anonymous):

A. 2x + 2y = 7 3x + 3y = 3 B. 2x + 2y = 7 5x + 5y = 10 C. 7x + 5y = 4 3x + 3y = 3 D. 7x + 5y = 4 5x + 5y = 10

OpenStudy (anonymous):

these are my only answers to choose from though?

OpenStudy (blurbendy):

when we added the first equation, x - y + 2z = -2, not 2, so when we added it the third equation, we should have gotten 4

OpenStudy (blurbendy):

have to be careful with the signs

OpenStudy (anonymous):

so which one is the answer then?

OpenStudy (anonymous):

cause im soo confused

OpenStudy (blurbendy):

our goal is to eliminate z. all we did to eliminate z for the first and third equations was first: multiply the third equation by 2, so we could get -2z to cancel with +2z in the first equation that gave us: x - y + 2z = -2 + 6x +6y -2z = 6 = 7x + 5y = 4

OpenStudy (blurbendy):

do you understand that part?

OpenStudy (anonymous):

okay .. get that part now.

OpenStudy (anonymous):

what do we do next?

OpenStudy (blurbendy):

now, to eliminate z in the second and third equations, all we have to do is add them, since the z's will cancel 2x + 2y + z = 7 + 3x + 3y – z = 3 = ???

OpenStudy (anonymous):

5x + 5y = 10

OpenStudy (blurbendy):

good job

OpenStudy (anonymous):

which means it would be D. 7x + 5y = 4 5x + 5y = 10

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (blurbendy):

yw